OpenStudy (anonymous):

Marion fills a bucket of water. The capacity of the bucket is 2 quarts. What is the capactiy of Marions bucket in gallons? 1 gallon=4 quarts

OpenStudy (anonymous):

1 gallon = 4 quarts x gallons = 2 quarts 1/x = 4/2 1/x = 2 x = 1/2 = 0.5 gallons

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw